OpenStudy (anonymous):

solve 4X^3-24X^2+44X-24 hai dam to answer do?

OpenStudy (campbell_st):

well here is part of the solution x = 1 is a zero \[p(1) = 4(1)^3 - 24(1)^1 + 44(1) - 24...so...P(1) = 0\] therefore (x -1) is a factor you can also take 4 out as a common factor so you have \[P(x) = 4(x^3 - 6x^2 + 11x - 6)\] so a little polynomial division  so you polynomial is now \[P(x) = 4(x -1)(x^2 - 5x + 6)\] just factor the quadratic hope this helps.